# 11email: {giovanni.catalani}[at]airbus.com

### v1

@@ 3 Experiments
We experimentally demonstrate the validity of the proposed approach on two steady state PDE benchmarks: (i) the hyper-elastic material dataset and (ii) the AirFRANS Dataset, consisting of numerical solutions of the incompressible Navier-Stokes equations on 2D airfoils. Additionally, we provide an experiment on learning more complex geometry representations of multi-element airfoils.The implementations and the code to reproduce the experiments can be found in the dedicated repository: https://anonymous.4open.science/r/enf2enf-322F.
3.1 The Hyper-elastic Material Dataset
3.1.1 Description
We consider a hyper-elastic material benchmark problem commonly used for geometry aware inference [23]. The boundary-value problem is defined on a unit cell domain, featuring an arbitrarily shaped void at its center. The underlying finite element solver employs approximately 100 quadrilateral elements. The input configuration is provided as a point cloud of around 1000 points representing the geometry, and the target output is the corresponding stress field. The dataset comprises 1000 training samples and 200 test samples. Although not representing a large scale PDE benchmark, this is a good demonstrative example that the proposed model has competitive performance beyond fluid dynamics applications. More details about the dataset are given in the Supplementary Materials.
3.1.2 Experimental Details
For the hyper-elastic material dataset, our approach employs an encoder-decoder architecture without concatenating any global features. For each input function, we compute a geometry descriptor by encoding the displacement field between the average mesh position and the current sample’s points, as done in previous works [32]. The obtained latent representation is defined on a point cloud of 9 nodes with 8-dimensional features, which is subsequently used to condition the output decoder model for the stress field. Baseline models such as FNO and UNet are limited to uniform grids: these methods rely on interpolating the unstructured input data onto a uniform grid, processing the data, and then re-interpolating the output back onto the original mesh. We also compare against DeepONet [26], functa2functa (CORAL with random Fourier features), and CORAL (which employs SIREN activation functions).
3.1.3 Results
Table 1 reports the L2 mean relative errors on the hyper-elastic material dataset. The results indicate that the performance of FNO and UNet is limited, likely due to the interpolation steps required to map between unstructured and uniform grids. Both CORAL and functa2functa achieve lower errors, with CORAL performing best among the baselines. Although the difference between local (enf2enf) and global latent encoding (functa2functa) is less pronounced for this relatively simpler PDE where multiscale information is not as critical, the proposed enf2enf still yields competitive results.
@@ 3.1.3 Results
Volume $\downarrow$ Surface $\downarrow$ $C_{L}$downarrow $\rho_{L}$uparrow
Simple MLP 0.0081 0.0200 0.2108 0.9932
@@
functa2functa* 2.19e-2
enf2enf (Ours) 1.88e-2
3.2 The AirFRANS dataset
3.2.1 Description
@@
The AirFRANS dataset [5] provides a challenging benchmark for surrogate modeling in computational fluid dynamics, simulating the incompressible Reynolds-averaged Navier-Stokes (RANS) equations over a wide range of airfoil geometries and flow conditions. The geometries are systematically obtained through variations of the NACA 4- and 5-digit series airfoils, resulting in a diverse design space. The implicit distance (or SDF) field is defined over the entire mesh for each simulation as a continuous descriptor of the airfoil shapes. The flow conditions cover incompressible regimes (Mach numbers < 0.3) with angles of attack ranging from -5° to 15°, representing typical subsonic flight conditions. The Reynolds number, ranging from 2 to 6 Millions is set within realistic flight ranges (for relatively small low speed aircrafts). The simulations are performed using OpenFOAM [19]. The dataset features a large scale PDE problem of industrial relevance, discretized on high resolution C-grid meshes with approximately 200,000 points per geometry, refined towards the airfoil surface to capture boundary layer effects. Each simulation takes around 25 minutes to achieve convergence using 16 CPU cores of an AMD Ryzen Processor. We focus on the full task as 800 samples are used for training and 200 test samples for model evaluations. More details about the Airfrans Dataset are given in the Supplementary Materials.
3.2.2 Experimental Setting and Baselines
We mainly evaluate the model performance on the volumetric and surfacic pressure field predictions and the resulting lift coefficients, as these are the primary quantities of interest for aerodynamic design. Our methodology (enf2enf) relies on a two step training strategy: (i) Encoding, the encoder ENF takes as input the coordinates ($x,y$) and fits the implicit distance fields using latent point clouds (9 points with 8-dimensional latent features). (ii) Decoding, the fitted input latent point clouds are concatenated with the freestream conditions ($V_{x},V_{y}$) and decoded to the output fields by a similar ENF architecture which operates in decode-only mode (latent features are not updated during training). More details about implementation and hyperparameters are provided in Supplementary Materials.
In order to show the effect of the added inductive bias on accuracy and generalization, we compare our model to a strong Neural Field baseline based on meta-learned functa [12] (referred to as functa2functa). This implements the CORAL [32], but using Fourier Features rather than SIREN networks used in the original paper (this would otherwise lead to poor convergence on this dataset). Additional comparison are performed with state-of-the-art architectures based on: Graph Neural Networks (including GraphSAGE [17], MeshGraphNet [28], and Graph U-Net [15]), Neural Operators (GINO [25]), and Transformer-based models (Transolver [40]).
3.2.3 Results
@@
In Table 1, it can be observed that enf2enf achieves the lowest errors across volumetric and surface pressure predictions as well as the lift coefficient estimation. The performance gap becomes more evident especially on the surface pressure, as more complex flow features are localized close to the boundary, compared to the rest of the domain where the flow can be considered quasi-potential. Graph Neural Network (GNN) baselines—such as GraphSAGE and Graph U-Net tend to perform better than MeshGraphNet for large scale PDEs thanks to the use of pooling and aggregation operations over extensive neighborhoods. Similarly to Graph U-Net, the proposed method, performs feature encoding on a latent set of nodes. However, Graph U-Net relies on complex pooling operations and aggregation steps to produce its latent representation, which is ultimately determined by the discretization of the input graph. Instead, enf2enf as a neural operator approach, yields representations that are independent of the specific input mesh instance, and can be decoded continuously in the spatial domain, thus being better suited to handle such tasks. To assess the impact of explicit local feature encoding on capturing relevant aerodynamic phenomena, we next compare with global latent encoding methods such as functa2functa in Figure 2. This aspect is especially crucial in fluid dynamics applications, where localized phenomena—such as the leading edge suction peak—play a critical role in determining aerodynamic performance. As depicted in Figure 2, the surface $C_{p}$ plot reveals that the maximum error at the leading edge is mitigated in the enf2enf predictions. Moreover, the isobaric lines outputs match more closely the high fidelity simulation with lower levels of noise (overfitting) and improved physical compliance. While Transolver employs a similar inductive bias to capture local effects, it lacks a more explicit mechanism for encoding geometric information, which may account for its relatively inferior performance. Finally, the t-SNE visualizations in Figure 3 further illustrate that enf2enf effectively learns and disentangles local geometric features, yielding a well structured latent space. The local latent codes computed at farfield and central nodes (left and center panels) exhibit clear clustering that correlates with key geometric parameters, such as airfoil thickness. Notably, the farfield latent feature (left panel) shows little sensitivity to thickness variations because its spatial position is distant from the airfoil, and therefore it does not contribute significantly to reconstructing the SDF field in regions where thickness is a critical factor. In contrast, the 64-dimensional array latent representation from functa2functa (right panel) is less structured, as it aggregates local features into a single global representation. Consequently, the localized encoding in enf2enf enables the decoder to focus on significant aerodynamic features while filtering out less relevant local variations.
@@
Furhtermore, we evaluate discretization convergence on the Airfrans dataset by training each model at lower mesh resolutions (5k, 10k, 32k) and then testing on the full-scale meshes consisting of approximately 200k points. For direct inference, the entire high resolution mesh is processed in a single forward pass. For iterative inference, the mesh is divided into multiple subsamples, each processed in a separate forward pass until all 200k nodes are covered. Our results show that graph based methods (GraphSAGE and GUNet) perform poorly under direct inference when the test resolution greatly exceeds the training resolution, although iterative inference can partially mitigate this at the cost of multiple runs. For instance, GraphSAGE at 32k training resolution requires roughly six passes to cover a 200k test mesh, significantly increasing total inference time. By contrast, Transolver can handle zero-shot super-resolution training at 5k and testing at 200k yet sees degraded accuracy when the resolution gap grows. enf2enf remains discretization invariant and shows no difference between direct and iterative inference; it achieves high accuracy even when trained at much lower resolutions, thus avoiding the cost of large scale training while maintaining excellent performance at full resolution. All models achieve a remarkable 4 to 5 order of magnitude speedup at inference compared to high fidelity CFD. More results and visualization are provided in the Supplementary Material section.
@@
Figure 2: Visualization of pressure field predictions on an airfoil geometry. Top row: predictions from enf2enf (left) and functa2functa (middle) compared to the ground truth (right). Bottom row: corresponding absolute errors for each model and surface pressure distribution.
enf far-field                       enf close                          functa
@@
Figure 4: Discretization convergence analysis on the AirFRANS dataset. Left: Direct inference performance vs training resolution Center: Iterative inference; note that ENF is omitted in this panel due to its coordinate-based design. Right: Inference time histogram for models trained at a resolution of 32k points.
3.3 Additional Experiment: Multi-element Airfoil
To further demonstrate the benefits of localized geometric encoding, we conducted additional experiments on shape embedding of multi-element airfoils, by generating a synthetic dataset consisting of 200 geometric variations derived from the NASA GA(W)-I airfoil with a fowler flap. The variations are created through systematic translation and rotation of the flap element while maintaining the main airfoil fixed, representing typical high-lift device operations in aircraft design. For each configuration, we computed the signed distance function (SDF) field and trained both enf and functa encoders for reconstruction accuracy, evaluating the models’ capacity to learn structured representations of multi-component geometries with relative positioning and orientation variations. Dramatic improvement in reconstruction accuracy (4 orders of magnitude) are obtained using the ENF encoder: while functa encoder treats each geometric variation as an entirely new shape to be encoded in its global latent space, ENF translation-equivariant architecture naturally handles the relative positioning of elements through its local feature encoding. This allows the model to effectively decompose the geometry into its constituent elements and capture their spatial relationships. The qualitative results in Figure 5 display how the ENF-based encoder accurately reconstructs both the main airfoil and flap geometries across different configurations, maintaining sharp features and precise relative positioning.
Figure 5: Shape reconstruction results on the multi-element airfoil geometry. For functa latent of shape 32 are used. For enf latents point clouds of 4 nodes and 8-dimensional node features are used. First row: high-lift configuration. Second row: high-lift and high-drag configuration.
4 Discussion and Limitations
Respecting the underlying physics is paramount in PDE modeling. Although approaches that incorporate explicit physics constraints in the loss function, such as Physics Informed Neural Networks (PINNs) [30], have shown promise, they remain limited by challenges in generalizing to complex shape variations, accommodating varying boundary conditions, and scaling beyond simple test cases. By contrast, restricting the search space to functions that inherently respect the underlying symmetries of the PDE operator provides an alternative pathway for incorporating physical constraints. In this work we propose a framework to learn steady state PDE surrogates using the recently introduced Equivariant Neural Fields [39], combining the geometric inductive bias typical of CNNs and GNNs with the discretization invariance of neural fields. These properties are beneficial for large scale PDE modeling, where training on full resolution meshes is computationally prohibitive, and for capturing the effects of geometric variability. Nevertheless, several limitations remain. First, the positions of the latent points are currently fixed a priori. Preliminary experiments aimed at learning these positions resulted in increased computational cost and training instability. Future work may explore adaptive strategies to allow latent points to self organize around more relevant spatial regions, thereby potentially enhancing the model’s capacity to capture critical geometric features. Additionally, while enf2enf has been demonstrated on two dimensional problems, its extension to three-dimensional applications remains to be investigated. In 3D, a larger number of latent points will likely be necessary to adequately cover the domain, which may introduce further challenges in terms of model complexity and computational efficiency.
@@ 5 Conclusion
In this work, we introduced a novel geometry aware operator learning framework that leverages Equivariant Neural Fields (ENFs) within an encoder–decoder architecture to address steady state PDEs on complex geometries. The method encodes the input geometry using localized latent embeddings and directly decodes these representations into continuous output fields, thereby achieving discretization invariance. Extensive experiments on a large scale aerodynamic dataset (AirFRANS) and a hyper-elastic material benchmark demonstrate that our approach attains competitive and in many cases superior performance compared to state-of-the-art baselines, including graph based methods, transformer models, and alternative neural field techniques.
{credits}
[…]
The authors have no competing interests.
Appendix 0.A Implementation Details
The implementation of our model architecture, training scripts, and dataset processing pipelines is available at https://anonymous.4open.science/r/enf2enf-322F. This repository contains all the code necessary to reproduce our experiments. Our proposed enf2enf model is implemented in JAX, building upon the original Equivariant Neural Fields codebase. For the baseline, we generally use the existing codebases in Pytorch.
0.A.1 enf2enf
@@ Appendix 0.A Implementation Details
0.A.1.1 AirFRANS
For the AirFRANS dataset, we initialize the latent positions in a bounding box $[-0.5,0.25]\times[-0.25,0.25]$ closer to the airfoil surface, as most of the interesting flow phenomena occur in this region. The latent feature dimension in the decoder is increased from 8 to 10 to accommodate the concatenation of inflow conditions (velocity components) that globally condition the output field generation. The dataset is normalized by standardizing the output pressure fields and scaling the input coordinates to $[-1,1]$ through min-max normalization. Following the ’full’ task setup, we use 800 samples for training and 200 for testing. Due to the large mesh sizes ($\sim 200,000$ points), we dynamically downsample to 10,000 points during both encoder and decoder training. At inference time, the continuous nature of Neural Fields allows us to evaluate on the full resolution meshes.
0.A.1.2 Elasticity
For the elasticity problem, latent positions are initialized in a larger bounding box $[-0.75,0.75]\times[-0.75,0.75]$ to better cover the unit cell domain. The inputs and outputs are normalized following the same strategy as AirFRANS. Given the relatively small mesh sizes ($\sim 1000$ points), we can process the entire point cloud during training without downsampling, leading to better reconstruction of sharp features near the void boundaries.
0.A.2 Baseline Models
@@
0.A.2.1 Neural Field Baselines
We implement functa2functa, based on the CORAL architecture but modified to use Random Fourier Feature encoding instead of the original SIREN networks. To ensure a fair comparison, we employ a 64-dimensional latent space, which approximately matches our model’s representational capacity (9 latent points × 8 features). The Fourier Feature encoding parameters are kept consistent with our approach to isolate the impact of the architectural differences.
0.A.2.2 Graph-Based Approaches
@@
0.A.2.3 Neural Operator Methods
@@
0.A.3 Hardware and Runtime
All experiments were performed on NVIDIA A100 GPUs with 40GB of memory. Training times and epochs for different models (only for the models retrained on our hardware) are reported in Table 3.
Table 3: Training Time and Epochs Across Different Models and Datasets
AirFRANS Elasticity
Model Time (hrs) Epochs Time (hrs) Epochs
enf2enf (encoder/decoder) 5.7/12.9 800/2500 4 1500
functa2functa (encoder/process/decoder) 15 2000 10 2000
GraphSAGE 5.8 400 3 1000
Graph U-Net 8.5 400 5 1000
Transolver 20.7 400 - -
Training times include all necessary pre-processing steps, such as mesh downsampling for graph-based methods and interpolation for grid-based approaches. For enf2enf and functa2functa, we report encoder and decoder training times separately as they are trained sequentially.
Appendix 0.B Datasets Details
0.B.1 AirFRANS Dataset

### v2

@@ 4 Experiments
We experimentally demonstrate the validity of the proposed approach on two steady state PDE benchmarks: (i) the hyper-elastic material dataset and (ii) the AirfRANS Dataset, consisting of numerical solutions of the incompressible Navier-Stokes equations on 2D airfoils. Additionally, we provide an experiment on learning more complex geometry representations of multi-element airfoils in the Appendix A. The implementations and the code to reproduce the experiments can be found in the dedicated repository: https://github.com/giovannicatalani/enf2enf_pytorch.
The Hyper-elastic Material Dataset
We consider a hyper-elastic material benchmark problem commonly used for geometry aware inference [13]. The boundary-value problem is defined on a unit cell domain, featuring an arbitrarily shaped void at its center. The underlying finite element solver employs approximately 100 quadrilateral elements. The input configuration is provided as a point cloud of around 1000 points representing the geometry, and the target output is the corresponding stress field. The dataset comprises 1000 training samples and 200 test samples. Although not representing a large scale PDE benchmark, this is a good demonstrative example that the proposed model has competitive performance beyond fluid dynamics applications. More details about the dataset are given in the Supplementary Materials. Table 1 reports the L2 mean relative errors on the hyper-elastic material dataset. The results indicate that the performance of FNO and UNet is limited, likely due to the interpolation steps required to map between unstructured and uniform grids. The two neural field based approaches, CORAL and the proposed enf2enf achieve the lowest relative errors, likely due to the explicit geometric encoding which is crucial for this task.
@@
Volume $\downarrow$ Surface $\downarrow$ $C_{L}$ $\downarrow$ $\rho_{L}$ $\uparrow$
@@
enf2enf 1.88e-2
The AirfRANS dataset
@@
The AirfRANS dataset [24] provides a challenging benchmark for surrogate modeling in computational fluid dynamics [25], simulating the incompressible Reynolds-averaged Navier-Stokes (RANS) equations over a wide range of airfoil geometries and flow conditions. The geometries are systematically obtained through variations of the NACA 4- and 5-digit series airfoils, resulting in a diverse design space. The implicit distance (or SDF) field is defined over the entire mesh for each simulation as a continuous descriptor of the airfoil shapes. The flow conditions cover incompressible regimes with angles of attack ranging from -5° to 15°, representing typical subsonic flight conditions. The dataset features a large scale PDE problem of industrial relevance, discretized on high resolution C-grid meshes with approximately 200,000 points per geometry, refined towards the airfoil surface to capture boundary layer effects. We focus on the full task as 800 samples are used for training and 200 test samples for model evaluations. More details about the AirfRANS Dataset are given in the Supplementary Materials. We mainly evaluate the model performance on the volumetric and surfacic pressure field predictions and the resulting lift coefficients, as these are the primary quantities of interest for aerodynamic design.
@@
In Table 1, it can be observed that enf2enf achieves the lowest errors across volumetric and surface pressure predictions as well as the lift coefficient estimation. The performance gap becomes more evident, especially on the surface pressure, as more complex flow features are localized close to the boundary, compared to the rest of the domain where the flow can be considered quasi-potential. Graph Neural Network (GNN) baselines such as GraphSAGE and Graph U-Net tend to perform better than MeshGraphNet for PDEs on large meshes thanks to the use of pooling and aggregation operations over extensive neighborhoods. Similarly to Graph U-Net, the proposed enf2enf, performs feature encoding on a latent set of nodes. However, Graph U-Net relies on complex pooling operations and aggregation steps to produce its latent representation, which is ultimately determined by the discretization of the input graph. Instead, enf2enf as a neural operator approach, yields representations that are independent of the specific input mesh instance, and can be decoded continuously in the spatial domain, thus being better suited to handle such tasks.
To assess the impact of explicit local feature encoding on capturing relevant aerodynamic phenomena, we next compare with global latent encoding methods such as CORAL in Figure 2. This aspect is especially crucial in fluid dynamics applications, where localized phenomena such as the leading edge suction peak play a critical role in determining aerodynamic performance. As depicted in Figure 2, the surface $C_{p}$ plot reveals that the maximum error at the leading edge is mitigated in the enf2enf predictions. Moreover, the isobaric lines outputs match more closely the high fidelity simulation with lower levels of noise (overfitting) and improved physical compliance.
Discretization Convergence
@@
Furthermore, we evaluate discretization convergence on the AirfRANS dataset by training each model at lower mesh resolutions (5k, 10k, 32k) and then testing on the full-scale meshes consisting of approximately 200k points. For direct inference, the entire high resolution mesh is processed in a single forward pass. For iterative inference, the mesh is divided into multiple subsamples, each processed in a separate forward pass until all 200k nodes are covered. Our results show that graph based methods (GraphSAGE and GUNet) perform poorly under direct inference when the test resolution greatly exceeds the training resolution, although iterative inference can partially mitigate this at the cost of multiple runs. For instance, GraphSAGE at 32k training resolution requires roughly six passes to cover a 200k test mesh, significantly increasing total inference time. By contrast, Transolver can handle zero-shot super-resolution training at 5k and testing at 200k, yet sees degraded accuracy when the resolution gap grows. enf2enf remains discretization invariant and shows no difference between direct and iterative inference; it achieves high accuracy even when trained at much lower resolutions, thus avoiding the cost of large scale training while maintaining excellent performance at full resolution. All models achieve a remarkable 4 to 5 orders of magnitude speedup at inference compared to high fidelity CFD.
@@
Figure 2: Visualization of pressure field predictions on an airfoil geometry. Top row: predictions from enf2enf (left) and CORAL (middle) compared to the ground truth (right). Bottom row: corresponding absolute errors for each model and surface pressure distribution.
@@
Figure 3: Discretization convergence analysis on the AirfRANS dataset. Left: Direct inference performance vs training resolution Center: Iterative inference; note that enf2enf is omitted in this panel due to its coordinate-based design. Right: Inference time histogram for models trained at a resolution of 32k points.
@@ 5 Conclusion
We introduced a geometry-aware operator learning framework leveraging spatially grounded neural fields within an encoder-decoder architecture for steady-state PDEs on complex geometries. Unlike approaches that incorporate explicit physics constraints in loss functions (e.g., Physics Informed Neural Networks (PINNs) [26]), our method restricts the search space to functions respecting underlying PDE symmetries through translation equivariance, providing an alternative pathway for incorporating physical constraints. Our approach leverages the discretization invariance of neural fields, beneficial for large-scale PDE modeling where training on full-resolution meshes is computationally prohibitive. Experimental validation on aerodynamic (AirfRANS) and material modeling benchmarks demonstrates competitive performance compared to graph-based methods, recent transformer models, and alternative neural field techniques, particularly for surface phenomena where local geometric features are critical. Extension to three-dimensional problems requires investigation, likely necessitating more latent points and introducing computational complexity challenges.
Aknowledgements
The experiments presented in this paper have been run on the ISAE-Supaero HPC cluster PANDO. This work was supported by Agence Nationale de la Recherche (ANRT), through CIFRE PhD Fellowship sponsored by Airbus Operations SAS and ISAE-Supaero.
Appendix A Additional Experiment: Multi-element Airfoil
To further demonstrate the benefits of localized geometric encoding, we conducted experiments on multi-element airfoil shape reconstruction. We generated 200 geometric variations of the NASA GA(W)-I airfoil with a Fowler flap by systematically translating and rotating the flap while keeping the main airfoil fixed, representing typical high-lift device operations. We compare two encoding approaches for SDF field reconstruction:
  • $\bullet$
    Global encoding (CORAL/Functa): Uses a 32-dimensional global latent vector.
  • $\bullet$
    Local encoding (enf2enf): Uses 4 spatially-anchored points with 8-dimensional features each.
Dramatic improvement in reconstruction accuracy (3 orders of magnitude) are obtained using the local encoder: while global encoding treats each geometric variation as an entirely new shape to be encoded in its global latent space, the translation-equivariant architecture naturally handles the relative positioning of elements through its local feature encoding. This allows the model to effectively decompose the geometry into its constituent elements and capture their spatial relationships. The qualitative results in Figure 4 display how the local encoder accurately reconstructs both the main airfoil and flap geometries across different configurations, maintaining sharp features and precise relative positioning.
Figure 4: Shape reconstruction results on the multi-element airfoil geometry. Red Line indicates the 0-level Signed Distance Function, corresponding to the airfoil shape. Global latents of shape 32 are used. Local latent point-clouds of 4 nodes and 8-dimensional node features are used. First row: high-lift configuration. Second row: high-lift and high-drag configuration.
Appendix B Implementation Details
The implementation of our model architecture, training scripts, and dataset processing pipelines is available at https://github.com/giovannicatalani/enf2enf_pytorch. This repository contains all the code necessary to reproduce our experiments. Our proposed enf2enf model is implemented in PyTorch. For the baseline, we generally use the existing codebases in Pytorch.
B.1 enf2enf
@@ Appendix B Implementation Details
B.1.1 AirfRANS experiments
For the AirfRANS dataset, we initialize the latent positions in a bounding box $0.5,0.25]\times[-0.25,0.25]$ closer to the airfoil surface, as most of the interesting flow phenomena occur in this region. The latent feature dimension in the decoder is increased from 8 to 10 to accommodate the concatenation of inflow conditions (velocity components) that globally condition the output field generation. The dataset is normalized by standardizing the output pressure fields and scaling the input coordinates to [1,1][-1,1] through min-max normalization. Following the ’full’ task setup, we use 800 samples for training and 200 for testing. Due to the large mesh sizes ($\sim 200,000$ points), we dynamically downsample to 10,000 points during both encoder and decoder training. At inference time, the continuous nature of Neural Fields allows us to evaluate on the full resolution meshes.
B.1.2 Elasticity experiments
For the elasticity problem, latent positions are initialized in a larger bounding box $0.75,0.75]\times[-0.75,0.75]$ to better cover the unit cell domain. The inputs and outputs are normalized following the same strategy as AirfRANS. Given the relatively small mesh sizes ($\sim 1000$ points), we can process the entire point cloud during training without downsampling, leading to better reconstruction of sharp features near the void boundaries. For each input function, we compute a geometry descriptor by encoding the displacement field between the average mesh position and the current sample’s points, as done in previous works [14].
B.2 Baseline Models
@@
Neural Field Baselines
We implement the CORAL architecture, modified to use Random Fourier Feature encoding instead of the original SIREN networks. To ensure a fair comparison, we employ a 64-dimensional latent space, which approximately matches our model’s representational capacity (9 latent points × 8 features). The Fourier Feature encoding parameters are kept consistent with our approach to isolate the impact of the architectural differences.
Graph-Based Approaches
@@
Neural Operator Methods
@@
Appendix C Datasets Details
C.1 AirfRANS Dataset